OpenStudy (anonymous):

William wants to reserve a room at his swim club for a birthday party. The room costs $500 for 6 hours. William must also pay a $125 deposit. What percent of the total cost of the room is the deposit ?? A. 4% B. 20% C. 25% D. 83% Please help i will give a medal and a fan!!!!please

OpenStudy (anonymous):

@chaise help

OpenStudy (anonymous):

The answer is a because if u multiply 500 by 6 and add the deposit it gives u 3125 then u cross multiply X over 100 and 125 over 3125. 100 times 125 is 12500 and divided by 3125 is 4.

OpenStudy (texaschic101):

deposit = 125 cost of room = 500 + 125 = 625 125/625 = 0.20 = 20%

OpenStudy (anonymous):

Who is right???

OpenStudy (texaschic101):

this might be wrong....I can't figure out if the total cost of the room is including the deposit or not ?

OpenStudy (anonymous):

So it's A???is it right??

OpenStudy (texaschic101):

that depends.. if the total cost of the room includes the deposit...then it is 20% but if the total cost of the room does not include the deposit, then it is 25%

OpenStudy (anonymous):

Because it says William must also pay $125

OpenStudy (texaschic101):

ok.....then the total cost of the room is 500 + 125 = 625 125/625 = 0.20 = 20%

OpenStudy (anonymous):

Thanks for helping...